# Beyond Natural Language: LLMs Leveraging Alternative Formats for Enhanced Reasoning and Communication

### v1

@@ A.1 Models
For OpenAI’s models, we use gpt-3.5-turbo-1106 and gpt-4-1106-preview. For Gemini pro, there is no clear official version, we use the Gemini pro in 2024.1.
@@ A.2 Dataset Pre-Processing
[…]
@@ Appendix B Prompts
We list the prompts we used in this work in Tables 7, 8, 10, 11, 9, 13 and 12.
@@
At the end of your response, you must give your answer in the form of "the answer is: no" or "the answer is: yes". Let’s think step-by-step.
AutoForm:
@@
[…]
CoT:
Solve the problem presented below:
\${task_description}
RESPONSE GUIDELINES:
1. Think step by step.
2. Concluding with the Answer (Optional): End your response with "Answer: {answer}", where {answer} is the final result of your problem-solving process. The {answer} should be a single capital letter.
3. You can use tools to execute Python code to help you do some calculations. You should not finish the whole problem in the code.
4. Only when you end your response with "Answer: {answer}" will the dialogue end.
AutoForm:
Solve the problem presented below:
\${task_description}
RESPONSE GUIDELINES:
1. Initial State Representation: Begin by providing a clear and detailed representation of the initial state or conditions of the problem. It could be code, pseudocode, JSON, markdown table, logical operators, or math equation.
3. Concluding with the Answer (Optional): End your response with "Answer: {answer}", where {answer} is the final result of your problem-solving process. The answer should be a single capital letter.
4. You can use tools to execute Python code to help you do some calculations. You should not finish the whole problem in the code.
5. Only when you end your response with "Answer: {answer}" will the dialogue end.
Tool:
  "name": "exec_python",
  "description": "Execute a Python script.",
  "parameters": {
    "type": "object",
    "properties": {
      "code": {
        "type": "string",
        "description": "The Python code to execute. The code should print the result to stdout.",
    "required": ["code"],
[…]
@@ Appendix D Examples
In this section, we provide several response examples for each LLM on each task, so that readers can have a more intuitive understanding of the effects after using AutoForm. Tables 70, 71, 75, 76, 72, 73 and 74 present the examples of using AutoForm, which corresponds to the results in Table 2. The examples for two-step AutoForm, which corresponds to Table 3, are presented in Tables 50, 51, 52, 53, 54, 55, 56, 57, 58, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 14, 15, 16, 17, 18, 19, 20, 21, 22, 41, 42, 43, 44, 45, 46, 47, 48 and 49. For the multi-agent communication experiment, we present the examples in Tables 59, 60, 61, 62, 63, 64, 65, 66, 67, 68 and 69.
@@ Conclusion:
**Step-by-Step Solution Process:**
1. Andraya does not flip the coin. This means that the state of the coin remains the same.
@@ 2. Step-by-Step Solution Process:
Given that the mantle clock was left on the edge of the counter by Ed and was found on the floor, it is likely that the intruder knocked it over, causing it to stop at 10:59 p.m. This suggests that the break-in occurred around this time.
@@ 7. Divide both sides by 19:
[…]
The discriminant in this case is (-5)${}^{2}$ 4*1*k = 25 - 4k. For the equation to have two real and distinct roots, 25 - 4k must be greater than 0. Solving for k, we get k < 25/4.
[…]

### v2

@@ A.1 Models
For OpenAI’s models, we use gpt-3.5-turbo-1106 and gpt-4-1106-preview. For Gemini pro, we use the Gemini pro 1.0 in 2024.1.
@@ A.2 Dataset Pre-Processing
For Logic Grid, Information Essentiality and Minute Mysteries QA that are from Big-Bench, we download the dataset from the official repo [^1] https://github.com/google/BIG-bench. For Coin Flip, we download the dataset from https://huggingface.co/datasets/skrishna/coin_flip, and use the first 500 examples in the test set. For AQuA, we download the dataset from https://huggingface.co/datasets/aqua_rat/ and use its test set.
@@ A.3 Metrics
Dataset Model Math Equation Unordered List Ordered List Markdown Table Multi-level List Logical Expression
Logic Grid GPT-3.5 50.0 47.0 46.0 48.5 45.0 41.0
GPT-4 58.0 65.0 59.0 55.0 62.0 59.0
Gemini Pro 42.0 50.5 48.0 49.0 48.0 49.0
Coin Flip GPT-3.5 71.0 39.0 86.0 66.0 19.2 48.0
GPT-4 75.0 95.0 100.0 99.0 98.0 71.4
Gemini Pro 47.0 61.6 60.4 63.4 56.4 59.6
Info Essen GPT-3.5 22.0 30.8 30.8 29.4 30.8 30.8
GPT-4 73.5 73.5 76.4 73.5 76.4 75.0
Gemini Pro 45.6 29.4 32.4 29.4 44.1 38.2
MM QA GPT-3.5 22.2 22.2 23.2 27.1 22.7 20.2
GPT-4 39.9 41.9 37.7 36.9 42.4 37.9
Gemini Pro 29.6 27.6 25.6 27.6 28.1 24.1
AQuA GPT-3.5 63.4 62.2 63.0 56.7 59.8 66.0
GPT-4 76.4 72.4 80.3 78.0 78.0 79.8
Gemini Pro 56.7 57.5 59.8 47.2 55.5 56.3
Table 6: Effectiveness of various formats across different models and datasets.
Hotpot QAseparate ${}_{\text{separate context}}$
Model RougeL # Tokens $\bm{\Delta}$Tokens
GPT-3.5 + GPT-3.5 0.62 369.3 -
+AutoForm 0.53 286.6 -22.4%
GPT-4 + GPT-4 0.68 151.0 -
+AutoForm 0.69 100.0 -33.8%
Table 7: Performance on HotpotQA under the separate context setting. The supporting facts are distributed between two agents, necessitating inter-agent communication for aggregating information and deriving the correct answer.
@@ Appendix B Prompts
We list the prompts we used in this work in Tables 9, 10, 12, 11, 14 and 13.
@@
At the end of your response, you must give your answer in the form of "the answer is: no" or "the answer is: yes". Let’s think step-by-step. AutoForm:
@@ Appendix C Additional Experimental Results for Multi-Agent Communication
Additionally, we conduct experiments with a specific setting for HotpotQA to evaluate the performance of LLMs when supporting facts are distributed between two agents. This setting, termed separate context, ensures that the supporting facts are divided and distributed between the agents, necessitating effective communication to aggregate the information and derive the correct answer.
The results presented in Table 7 reveal significant insights into the performance of LLMs under the separate context setting. The application of the AutoForm mechanism shows a distinct impact on both models. For GPT-3.5, the RougeL score slightly decreases from 0.62 to 0.53, with a corresponding 22.4% reduction in the number of tokens generated. In contrast, the GPT-4 models exhibit an increase in performance with the AutoForm mechanism, achieving a RougeL score of 0.69 and a 33.8% reduction in token usage. This improvement highlights the effectiveness of AutoForm in not only maintaining but enhancing the performance of more advanced models while optimizing resource utilization.
@@ Appendix D Examples
In this section, we provide several response examples for each LLM on each task, so that readers can have a more intuitive understanding of the effects after using AutoForm. Tables 71, 72, 76, 77, 73, 74 and 75 present the examples of using AutoForm, which corresponds to the results in Table 2. The examples for two-step AutoForm, which corresponds to Table 3, are presented in Tables 51, 52, 53, 54, 55, 56, 57, 58, 59, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 15, 16, 17, 18, 19, 20, 21, 22, 23, 42, 43, 44, 45, 46, 47, 48, 49 and 50. For the multi-agent communication experiment, we present the examples in Tables 60, 61, 62, 63, 64, 65, 66, 67, 68, 69 and 70.
@@ Conclusion:
**Step-by-Step Solution Process:** 1. Andraya does not flip the coin. This means that the state of the coin remains the same.
@@ 2. Step-by-Step Solution Process:
- …… Given that the mantle clock was left on the edge of the counter by Ed and was found on the floor, it is likely that the intruder knocked it over, causing it to stop at 10:59 p.m. This suggests that the break-in occurred around this time.
@@ 7. Divide both sides by 19:
To find the largest integral value of ’k’ for which the quadratic equation x2 - 5x + k = 0 will have two real and distinct roots, we can use the discriminant formula. The discriminant (D) of the quadratic equation ax2 + bx + c = 0 is given by D = b2 - 4ac. For the equation x2 - 5x + k = 0, the discriminant must be greater than 0 for the equation to have two real and distinct roots.
The discriminant in this case is (-5)2 - 4*1*k = 25 - 4k. For the equation to have two real and distinct roots, 25 - 4k must be greater than 0. Solving for k, we get k < 25/4.
The largest integral value of k that satisfies this condition is 6. Therefore, the largest integral value of k for which the quadratic equation x2 - 5x + k = 0 will have two real and distinct roots is 6.